Phillips' DEIMOS News and Progress

  • DEIMOS Has First Light June 3 (04 Jun)
  • The Hawaii Report (31 May)
  • Red Mosaic Assembly (20 Mar)
  • DEIMOS Structure Ships Feb 7 (19 feb)
  • Optics: Camera (21 Mar)
  • Dewar Window AR Coat

    Click on pictures for larger versions.

    DEIMOS Has First Light June 3

    First light went almost flawlessly. The weather was beautiful, we were integrated with the telescope and guider within about 2 hours after opening up. When we pointed to the first star (in twilight), it fell near the center of the guider! By the end of the night we had aligned with a slitmask to check the astrometry and had acquired all the basic photometric and spectrophotometric calibration data. Pictures soon!

    News from Hawaii

    11 May: Drew has been in Hawaii for the past week. We are currently in the process of adjusting the kinematic mounts to place DEIMOS's rotation axis on the elevation axis (ie optical axis) of the telescope. This has proceeded in 3 stages: (a) CARA put in a cross-hair on one elevation axis, and a "target" on the AO optical rotator, which is believed to be precisely on axis; (b) DEIMOS moved into observing position, and autocollimator aligned on the two points above defining the elevation axis; and (c) DEIMOS aligned to the autocollimator. Due to the small sight angle through DEIMOS, we have had to perform (b) and (c) twice, first to place DEIMOS approximately on-axis so that we had a clear sight of the cross-hair/target defining points, and then for the final alignment of autocollimator and instrument. We are in the process of finishing with (c) for the final pass.

    30 May: Drew is back in Hawaii. This page needs a lot of work, but for now here are some pictures of current problems:

    Bubbles in Optical Couplant fluid between Elem 5 and 6:

    click for enlargement

    Another photo with Flash lighting

    Degradation of Collimator and Tent mirror Coatings:

    Collimator coating detail 2 The largest spot is several mm across including outer halo.

    31 May: PRM light leak appears to be fixed. This was a problem where the PRM sensor, located on the grating box, seemed not to turn its LED completely off, leading to a high background light level in long exposures. After we identified the PRM sensor LED as the culprit, Bob Kibrick found in Santa Cruz that very low current levels could indeed cause these LEDs to glow faintly. A fix was derived; the requested resistor installed in the interconnect box, and the latest "dark" exposures show no trace of the characteristic shadows cast by the PRM sensor LED.

    The PXL guide camera has been restored to health, images obtained and analyzed in preparation for first light on June 3.

    Other news: (To be added)

    Red Mosaic Assembly and Testing

    12 Feb: The red mosaic is assembled and scheduled to go into the Dewar shortly. Modifications have been made to the Dewar front plate to allow more room for the FCS devices, which provide the limiting tolerance. The Back-Focal-Distance (proper distance behind the Dewar window) is believed to be understood now.

    Examination of the so-called blue mosaic reveals a lot of contamination of two sorts:

  • tiny metal particals (copper possibly generated when the "cold-finger" is attached, and others (?) -- now being analyzed); and
  • apparent residue left by liquid water. It is speculated the O-ring may leak slightly, allowing water vapour to freeze out on the CCDs; upon warming, the layer of ice/frost collects in droplets.
    NB: The contamination is under investigation and the possible causes are mere speculation at this point. However, beyond being very careful with the reassembly and applying a suction force while assembling, there is no plan for modification or expected delay.

    14 Feb: The red mosaic was put in the Dewar yesterday, but it was dicovered that it could not be rotated sufficiently for proper alignment. New mount points are being machined to correct this.

    15 Feb: New mount points are manufactured and being fine-tuned. However, an illness is delaying final installation until next week.

    26 Feb: Red Mosaic has been rotated and levelled. One amplifier out of 16 is NOT functioning -- this is currently under investigation.

    28 Feb: The "non-functioning" amplifier appears to be resolved -- the CCD has some very hot pixels in the serial register near the amplifier which was causing alarming behavior, but upon cooling things look OK. Thus the Red Mosaic appears to be fully-functioning! Characterizations continue ...

    20 Mar: Voltages have been adjusted to bring linearity within the specified range on all 16 amplifiers. Amplifier 5A continues to show infrequent, intermittent problems. Fringing images at 9123A will be acquired tonight.

    DEIMOS Structure Ships Feb. 7

    DEIMOS lifted off regular truck The structure (barrel and cradle, in large white "gift box") was transfered to a "lowboy" trailer ...
    DEIMOS structure leaving campus ... and headed for the Port of Oakland early this morning, Feb. 7. A hardy group (photo by Gary Dorst) braved the drizzle to see it off.

    12 Feb: The latest word is that DEIMOS arrived safely at the port, was safely loaded and sailed on schedule.

    12 Feb: The second shipment to go by sea (the container) left Tuesday, Feb. 12, the scheduled day. The next scheduled shipment is an air shipment containing electronics, scheduled for next Tuesday. (Shipping Schedule.)

    14 Feb: DEIMOS structure is in Hawaii; see Heather Mietz's updates on the LOEN DEIMOS Homepage for additional details.

    19 Feb: DEIMOS structure is in the Keck II dome! See link above for more details/photos.

    Optics: Camera

    The camera has been disassembled into segments, and various tests and measurements are being performed. Goals here are to verify assembly and camera characteristcs and to fix sources of flexure in the camera. It was known beforehand that Element 3 had some radial motion -- this has been confirmed and extra shimming added. Some other problems have come to light (see below).

    The plan is to reassemble the camera shortly and perform both the COHU "double-pass" tests and some "Pinhole shadow tests" to measure the [hopefully-improved] level of image motion.

    19 Feb: The camera has been reassmbled; image motion has been reduced by a factor of 3! This is comparable to the measured flexure in the detector, which has the opposite sign, so we are satisfied with the current flexure level in the ecamera.

    26 Feb: There has been some investigation of asymmetric image quality. The source of this is not known -- it's hypothesized that the problem lies in elements 7 and 8, whose aspheric surfaces are known to be slightly decentered, and may have been assembled so that the decenter effects are adding rather than subtracting. However, ZEMAX modelling shows that the effect can be removed by appropriately decentering Group 4 (Elements 8 + 9), and doing this indeed balanced out the PSFs. The camera is ready for disassembly and packing, approximately on schedule.

    01 Mar The asymmetry discussed above has been tracked to a tilt of Group 4 wrt the optical axis.

    15 Mar The camera has been shipped and reassembled in Hawaii. Initial comparisons show that the comatic tails on the images have shrunk significantly by a temperature of 55F.

    20 Mar Cold camera images were obtained after applying the decenter to Group 4 needed to compensate for its tilt (01 Mar). Also, the spacing of Elem 9 with respect to Elem 8 was varied. Results show good images especially with the 60 mil shim. (Also, see version with saturated images.)

    21 Mar Archival COHU images were obtained; see mosaic at 0, 5 and 9 deg off axis. Note that the aliasing and elongation in X is an artifact of the COHU camera, as shown by rotating it while imaging the same PSF.


    The central element in the Group 3 triplet is a very-convex CaF2 lens, which developed a fracture during fabrication. The fracture, visible to the right in the photo, was examined and no changes were noticed compared to the time of original assembly.

    Group 3 was found to exhibit some gravity-dependent "wedge" due to a gradient in fluid pressure from the couplant fluid; this wedge contributes to image motion. The gradient (top-to-bottom) was measured at +/- 2.6 thou. The last element in Group 3 was found to have come off its axial defining points; this has been fixed, which reduced the wedge slightly, to +/- 2.0 thou.

    Group 1 (a doublet) also shows this wedge (at a level of +/- 1.5 thou), and it is believed Group 4 (also a doublet) would also have this problem, although not as badly.

    The wedge is probably responsible for the difference in image motion seen in the double-pass COHU tests compared to the "pinhole-shadow" tests (which would be largely insensitive to the effect of wedge). Furthermore, the measurements of the wedge seem to have a significant drift over timescales of minutes, which makes this a good candidate for the source of "image droop".

    14 Feb: Inputing the measured wedge into optical models shows that it precisely accounts for the "missing" image motion seen in the COHU tests. We now believe we understand all significant sources of flexure in the camera. It has also been discovered that the Delrin retaining ring on Element 2 (Group 1) would deform under normal loading; this ring was replaced with a stainless steel ring, which reduced the measured wedge in this group by half.

    Group 1 on turntable For completeness, here's a photo of Group 1 ...
    Elem.3 in Barrel ... Element 3 ("Group 2"), a single lens, the largest piece of CaF2 ever in an astronomical instrument ...
    Group 4 (front) ... and Group 4. This group moves axially to maintain the plate scale with temperature changes; the "bellows" mount (stainless-steel diagonal vanes) can be seen in spaces between the black flocking.
    Elem.9 in cell Element 9 (the field flattener) being recentered in its cell ...
    Dave Hilyard centering Elem.9 ... and relative to Group 4.

    Dewar Window AR Coat

    Front of Dewar, disassembly Jan02 When the dewar was removed from DEIMOS in early January, it was found that condensation on the dewar window has caused the $10,000 AR coating to "crizzle". We had incorrectly identified this pattern of cracks as frozen condensation.
    Dewar Window after Cleaning After wiping, the "crizzled" coating fell off. However, the AR coat is intact over most of the window. The current plan is to live with this as is.

    Last modified: 22 mar 02

    Andrew C. Phillips / Lick Observatory